# Quantitative polynomial cohomology and applications to $\textrm{L}^{p}$-measure equivalence

### v1

@@ 5.1 Filling functions of non-cocompact lattices in rank 1
Leuzinger [leuzinger-rank1] (elaborating on [young-filling-nilpotent]) and Gruber [gruber-filling] obtained bounds for filling functions of non-cocompact lattices of rank 1 Lie groups. These bounds, combined with Theorem 5.1, will allow us to deduce information about their unitary and Banach cohomology. We record that for the quaternionic hyperbolic space of dimension $4n$, we are not aware of any computations of filling functions in degrees above $n$ nor for the octonionic hyperbolic plane.
@@
$\bullet$ [leuzinger-rank1, Theorem 4] For $\mathbb{K}=\mathbb{R}$, we have:
@@
$\bullet$ [leuzinger-rank1, Theorem 5] For $\mathbb{K}=\mathbb{C}$, we have:
@@
$\bullet$ [gruber-filling, 5.9] For $\mathbb{K}=\mathbb{H}$, we have:
@@ 5.2 Consequences for unitary representations and $L^{p}$-cohomology
Let $\Lambda$ be a non-cocompact lattice in $G=\mathrm{Isom}_{0}(\mathbb{H}^{n}_{\mathbb{K}})$, where $\mathbb{H}^{n}_{\mathbb{K}}$ denotes the hyperbolic space of dimension $n$ over $\mathbb{K}=\mathbb{R},\mathbb{C}$ or the quaternions $\mathbb{H}$. Let $(\pi,\mathcal{H})$ be a unitary representation of $\Lambda$. Then if either:
@@
Let $d=\dim_{\mathbb{R}}\mathbb{K}$. In order to use Theorem E for $p=2$, we need that $2N_{k}<d(n+1)-2$. With our previous computation of $N_{k}$, this condition boils down to 2k<d(n+1)22k<d(n+1)-2 for $1\leq k<n$.
For $d=1$, this is only satisfied for $1\leq k<\frac{n-1}{2}$.
For $d=2,4$ this is satisfied as soon as $1\leq k<n$.
@@
The range of $k$ for which this result holds should be larger in the quaternionic case. The reason why we cannot go higher is because we do not know any estimates on the filling functions of lattices in $\mathbb{H}_{\mathbb{H}}^{n}$ in degrees between $n$ and $4n$. At least looking at the real and the complex case, we expect to reach all degrees below half of the dimension.
One particular Banach representation that is of particular importance is the regular representation in $L^{p}$. Its (continuous) cohomology is what we commonly call (continuous) group $L^{p}$-cohomology. Some properties that make $L^{p}$-cohomology interesting are the following.
$\bullet$ It is invariant by quasi-isometries [pansu95], [bourdon-remy-vanishings, 1.1].
$\bullet$ It has natural geometric interpretations: there are simplicial and de Rham versions of $L^{p}$-cohomology, which coincide with the group-theoretic version for well-behaved simplicial complexes and manifolds admitting nice geometric actions [bourdon-remy-vanishings, 3.2], [bourdon-remy-non-vanishing, A.1].
Pansu showed a vanishing result for de Rham $L^{p}$-cohomology of manifolds with pinched strictly negative sectional curvature [pansu08, Théorème A]. This applies in particular to non-compact globally symmetric spaces of rank 1, and by extension to their isometry groups and their cocompact lattices. As an application of Theorem 5.1, we obtain that non-cocompact lattices satisfy the same vanishing result as their ambient simple Lie groups of rank 1.
@@
$\bullet$ $\mathbb{K}=\mathbb{H}$, $1\leq k\leq n-1$ and $1\leq p<\frac{4n+k-1}{2k}$.
@@
But the induced representation $I^{p}(\ell^{p}(\Lambda)))$ acting on $L^{p}(X,\ell^{p}(\Lambda))\simeq L^{p}(G)$ is continuously equivalent to the regular representation of $G$ on $L^{p}(G)$. Cohomology of the latter representation is in turn equivalent to the de Rham $L^{p}$-cohomology of $\mathbb{H}^{n}_{\mathbb{K}}$. To sum up
@@
In [pansu08, Théorème A], Pansu showed that
@@
where $\delta$ is the pinching of the sectional curvature of $\mathbb{H}^{n}_{\mathbb{K}}$. For $\mathbb{K}=\mathbb{R}$, we have $\delta=-1$ and for $\mathbb{K}=\mathbb{C},\mathbb{H}$ we have $\delta=-\frac{1}{4}$. Thus, we obtain
@@
Using our computation of $N_{k}$, case by base verification yields that
@@
Antonio López Neumann
Université Paris Cité, Sorbonne Université, CNRS, IMJ-PRG, F-75013 Paris, France
alopez@impan.pl / lopezneumann@imj-prg.fr
Juan Paucar Zanabria
Université Paris-Cité, IMJ-PRG, Paris
jpaucar@imj-prg.fr

### v2

@@ 5.1 Filling functions of non-cocompact lattices in rank 1
Leuzinger [32] (elaborating on [47]) and Gruber [27] obtained sharp bounds for filling functions of non-cocompact lattices of rank 1 Lie groups. These bounds, combined with Theorem 5.1, will allow us to deduce information about their unitary and Banach cohomology. We are not aware of any precise computations of filling functions for lattices of the quaternionic hyperbolic space of dimension $4n$ in degrees above $n$ nor for the octonionic hyperbolic plane, nevertheless Wenger [46, 7.3] still provides some (likely non sharp) upper bounds for filling functions of Carnot groups which imply upper bounds for filling functions of all lattices in rank 1 simple Lie groups.
@@
$\bullet$ [32, Theorem 3 and 4] For $\mathbb{K}=\mathbb{R}$, we have:
@@
$\bullet$ [32, Theorems 3 and 5] For $\mathbb{K}=\mathbb{C}$, we have:
@@
$\bullet$ [32, Theorem 3] and [27, 5.9] For $\mathbb{K}=\mathbb{H}$, we have:
@@
We now turn to deriving some general upper bounds for the remaining cases.
Proposition 5.3.
Let $\Lambda$ be a non-cocompact lattice in $G=\mathrm{Isom}_{0}(\mathbb{H}^{n}_{\mathbb{K}})$, where $\mathbb{H}^{n}_{\mathbb{K}}$ denotes the hyperbolic space of dimension $n\geq 2$ over $\mathbb{K}=\mathbb{R},\mathbb{C}$, the quaternions $\mathbb{H}$ or the octonions $\mathbb{O}$ (in which case $n=2$). For $2\leq k\leq n-1$, we have:
$\mathrm{cFV}^{k}_{\Lambda}(t)\lesssim t^{\frac{2^{k}-1}{2^{k-1}-1}}.$
Proof.
Let $G=KAN$ denote the Iwasawa decomposition of $G$, where $K$ is a maximal compact subgroup of $G$, AA a maximal $\mathbb{R}$-split torus in $G$ and $N$ is the unipotent radical of the minimal parabolic subgroup containing $A$. By the Federer-Fleming deformation technique, the combinatorial filling functions $\mathrm{cFV}_{\Lambda}^{k}$ are equivalent to their more classical singular homological variants $\mathrm{FV}_{\Lambda}^{k}$ [5, Theorem 6.6]. The group $\Lambda$ acts properly and cocompactly on the space $X_{0}$ obtained by removing a $\Lambda$-invariant family of horoballs to the hyperbolic space $\mathbb{H}_{\mathbb{K}}^{n}=AN$, where the boundary of each horoball (horospheres) is isometric to the nilpotent Lie group $N$. As in the proof of [32, Theorem 5], we obtain that
$\mathrm{FV}_{\Lambda}^{k}\sim\mathrm{FV}_{N}^{k}$
for every $k\geq 2$. The Lie group $N$ is Carnot and at most 2-step nilpotent. Hence [46, 7.3] shows that for every $k\geq 2$ we have
$\mathrm{FV}^{k}_{N}(t)\lesssim t^{1+\frac{2^{k-1}}{1+2+\ldots+2^{k-2}}}=t^{\frac{1+2+\ldots+2^{k-1}}{1+2+\ldots+2^{k-2}}}=t^{\frac{2^{k}-1}{2^{k-1}-1}}.$
which implies the desired result. ∎
This result implies that we have the following general upper bound on the constant $N_{k}$:
$N_{k}\leq\prod_{j=2}^{k}\frac{2^{j}-1}{2^{j-1}-1}=2^{k}-1.$
@@ 5.2 Consequences for unitary representations and $L^{p}$-cohomology
Let $\Lambda$ be a non-cocompact lattice in $G=\mathrm{Isom}_{0}(\mathbb{H}^{n}_{\mathbb{K}})$, where $\mathbb{H}^{n}_{\mathbb{K}}$ denotes the hyperbolic space of dimension $n\geq 2$ over $\mathbb{K}=\mathbb{R},\mathbb{C}$, the quaternions $\mathbb{H}$ or the octonions $\mathbb{O}$. Let $(\pi,\mathcal{H})$ be a unitary representation of $\Lambda$. Then if one of the following holds:
@@
Let $d=\dim_{\mathbb{R}}\mathbb{K}$. In order to use Theorem E for $p=2$, we need that $2N_{k}<d(n+1)-2$. With our previous computation of $N_{k}$ obtained after stating Theorem 5.2, this condition reduces to 2k<d(n+1)22k<d(n+1)-2 for $1\leq k<n$.
For $\mathbb{K}=\mathbb{R},d=1$, this is only satisfied for $1\leq k<\frac{n-1}{2}$.
For $\mathbb{K}=\mathbb{C},d=2\text{ and }\mathbb{K}=\mathbb{H},d=4$ this is satisfied as soon as $1\leq k<n$.
For $\mathbb{K}=\mathbb{O},d=8$ and $n=2$, we use the inequality $N_{k}\leq 2^{k}-1$ obtained after Proposition 5.3 and see that the condition $2N_{k}<d(n+1)-2=22$ is satisfied for $k=1,2,3$.
@@
The range of $k$ for which this result holds should be larger in the quaternionic case. The reason we cannot go higher is because we do not know good estimates on the filling functions of (non-uniform) lattices in $\mathbb{H}_{\mathbb{H}}^{n}$ in degrees between $n$ and $3n$. At least looking at the real and the complex case, we expect to reach all degrees below half of the dimension.
In [5, Appendix A], it is shown that the real simple Lie group $G=\mathrm{Isom}_{0}(\mathbb{H}_{\mathbb{O}}^{2})$ has property $T_{3}] [$5, Definition 1.1]. The previous result shows that this property is inherited by its lattices.
Corollary 5.5.
Let $\Lambda$ be a non-cocompact lattice in $G=\mathrm{Isom}_{0}(\mathbb{H}^{2}_{\mathbb{O}})$. Then $\Lambda$ has property $T_{3}]$, that is, for every unitary representation $\pi$ of $\Lambda$, we have $H^{k}(\Lambda,\pi)=0$ for $k=1,2,3$.
An important Banach representation is the regular representation in $L^{p}$. Given a locally compact second countable group $G$, we denote by $H^{*}_{\mathrm{ct}}(G,L^{p}(G))$ the (continuous) cohomology of $G$ with values in the right regular representation on $L^{p}(G):=L^{p}(G,m_{G})$, where $m_{G}$ is a left Haar measure, we commonly call this space the (continuous) group $L^{p}$-cohomology of $G$. Some properties that make $L^{p}$-cohomology interesting are the following.
$\bullet$ It is invariant by quasi-isometries [37], [8, 1.1].
$\bullet$ It has natural geometric interpretations: there are simplicial and de Rham versions of $L^{p}$-cohomology, which coincide with the group-theoretic version for well-behaved simplicial complexes and manifolds admitting nice geometric actions [8, 3.2], [9, A.1].
Pansu showed a vanishing result for de Rham $L^{p}$-cohomology of manifolds with pinched strictly negative sectional curvature [38, Théorème A]. This applies in particular to non-compact globally symmetric spaces of rank 1, and by extension to their isometry groups and their cocompact lattices. As an application of Theorem 5.1, we obtain that non-cocompact lattices satisfy the same vanishing result as their ambient simple Lie groups of rank 1.
@@
$\bullet$ $\mathbb{K}=\mathbb{H}$, $1\leq k\leq n-1$ and $1\leq p<\frac{4n+k-1}{2k}$,
$\bullet$ $\mathbb{K}=\mathbb{O}$, $k=1,2,3$ and $1\leq p<\frac{15+k}{2k}$.
@@
But the induced representation $I^{p}(\ell^{p}(\Lambda))$ acting on $L^{p}(X,\ell^{p}(\Lambda))\simeq L^{p}(G)$ is continuously equivalent to the regular representation of $G$ on $L^{p}(G)$. Cohomology of the latter representation is in turn equivalent to the de Rham $L^{p}$-cohomology of $\mathbb{H}^{n}_{\mathbb{K}}$. To sum up
@@
In [38, Théorème A], Pansu showed that
@@
where $\delta$ is the pinching of the sectional curvature of $\mathbb{H}^{n}_{\mathbb{K}}$. For $\mathbb{K}=\mathbb{R}$, we have $\delta=-1$ and for $\mathbb{K}=\mathbb{C},\mathbb{H},\mathbb{O}$ we have $\delta=-\frac{1}{4}$. Thus, we obtain
@@
Using our computation of $N_{k}$, case by case verification yields that